OpenStudy (anonymous):

A slide 4.1 meters long makes an angle of 35º with the ground. To the nearest tenth of a meter, how far above the ground is the top of the slide?

OpenStudy (anonymous):

construct a simple right angled triangle with a hypotenuse of 4.1 metres making an angle of 35 degrees to the ground. using trigonometric ratios (in this case sin as we want the length of the side opposite to the angle) you arrive at sin(35)=height/4.1; rearranging gives you height=4.1*sin(35)

OpenStudy (anonymous):

which would give me 23.5?

OpenStudy (anonymous):

no it gives you 2.4 to the nearest tenth you've plugged 41 instead of 4.1 into you calculator

OpenStudy (anonymous):

1. Wayne used the diagram to compute the distance from Ferris, to Dunlap, to Butte. How much shorter is the distance directly from Ferris to Butte than the distance Wayne found? (1 point) (0 pts) 20 mi (0 pts) 25 mi (1 pt) 10 mi (0 pts) 35 mi 1 /1 point 2. A triangle has side lengths of 18 cm, 80 cm, and 81 cm. Classify it as acute, obtuse, or right. (1 point) (0 pts) obtuse (0 pts) right (1 pt) acute 1 /1 point 3. The area of a square garden is 200 m². How long is the diagonal? (1 point) (0 pts) 100 m (1 pt) 20 m (0 pts) m (0 pts) 400 m 1 /1 point 4. A conveyor belt carries supplies from the first floor to the second floor, which is 12 feet higher. The belt makes a 60° angle with the ground. How far do the supplies travel from one end of the conveyor belt to the other? Round your answer to the nearest foot. (1 point) (0 pts) 21 ft (0 pts) 7 ft (1 pt) 14 ft (0 pts) 17 ft 0 /1 point 5. Find the missing value to the nearest hundredth. tan ___ = 73 (1 point) (0 pts) 81.61° (0 pts) 64.61° (0 pts) 60.61° (1 pt) 89.22° 1 /1 point 6. Write the ratios for sin A and cos A. The triangle is not drawn to scale. (1 point) (0 pts) sin A = , cos A = (1 pt) sin A = , cos A = (0 pts) sin A = , cos A = (0 pts) sin A = , cos A = 0 /1 point